OpenStudy (theriverspirit):

How many grams of CO are needed to react with an excess of Fe2O3 to produce 225.5 g Fe? Reaction's equation: Fe2O3 (g) +3 CO (g) -- > 3 CO2(g) + 2 Fe(s)

OpenStudy (theriverspirit):

i need the molar mass of CO and Fe also I need the mass ratio and the correct answer

OpenStudy (aaronq):

Start by converting the mass of Fe to moles.

OpenStudy (theriverspirit):

i did its 4.037 moles

OpenStudy (aaronq):

Okay, cool. Now you use a ratio of the moles of the two things of interest (Fe and CO in this case) and the stoichiometric coefficients: \(\sf \dfrac{moles~of~Fe}{Fe's ~coefficient}=\dfrac{moles~of~CO}{CO's~coefficient}\) plug things in and solve for moles of CO

OpenStudy (theriverspirit):

so that would be 4.037/2=moles of CO/3?

OpenStudy (aaronq):

yes

OpenStudy (theriverspirit):

so the moles of CO is 6.055 Thank you and the ratio is 2.018/1?

OpenStudy (aaronq):

yeah, the question is asking for mass though. you still need to convert. a ratio of what?

OpenStudy (theriverspirit):

I typed the question wrong i needed to find the mole ratio which I did

OpenStudy (aaronq):

the mole ratio is the ratio of the coefficients you used

OpenStudy (theriverspirit):

which is 2:3 and the answer is 169.6 grams of CO

OpenStudy (aaronq):

cool cool

OpenStudy (theriverspirit):

okay and can we do one last problem that handles percent yields?